.epsi .eps .pspgflastimage

Skeleton is a shape descriptor which summarizes the general form of objects. 3D objects can be represented by their centerlines or their medial surface. In digital pictures the continous skeleton is difficult to determine, but there are several approach to estimate the skeleton in digital spaces.